OpenStudy (anonymous):

find the limit as x approaches 0 of sin 9x/x

OpenStudy (blockcolder):

\[\lim_{x\to0}\frac{\sin{9x}}{x}=\lim_{x\to0}\frac{\sin{9x}}{9x}\cdot 9=9\lim_{x\to0}\frac{\sin{9x}}{9x}=9\lim_{u\to0}\frac{\sin{u}}{u}\]